OpenStudy (anonymous):

For 2x – 3y = –7, when x = 4, What is the value of y?

OpenStudy (aravindg):

put x=4

OpenStudy (aravindg):

2x – 3y = –7 2*4-3y=-7 8-3y=-7

OpenStudy (aravindg):

can you do it from there ?

OpenStudy (texaschic101):

just sub the 4 in for x and solve for y

OpenStudy (anonymous):

@AravindG no.

OpenStudy (aravindg):

8-3y=-7 subtract 8 on both sides

OpenStudy (anonymous):

ok.

OpenStudy (aravindg):

What do you get ?

OpenStudy (anonymous):

1

OpenStudy (aravindg):

try again ?

OpenStudy (aravindg):

8-3y-8=-7-8

OpenStudy (anonymous):

so is it 5 or -5

OpenStudy (aravindg):

-3y=15 y=5 :)

OpenStudy (aravindg):

click best response if this helped

OpenStudy (aravindg):

-3y=-15 y=5 ^^

OpenStudy (anonymous):

Yay!

OpenStudy (aravindg):

:)